# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ B.3. Weak Isomorphisms
Let $\mathbb{Y}=(Y,\omega_{Y},\mu_{Y})=(\mathbf{Y},\mathrm{q})$ and $\mathbb{Y}^{\prime}=(Y^{\prime},\omega_{Y^{\prime}},\mu_{Y^{\prime}})=(\mathbf{Y}^{\prime},\mathrm{q}^{\prime})$ be finite networks such that $\mathbb{Y}\sim^{w}\mathbb{Y}^{\prime}$ (i.e., both belong to the same class $[\mathbb{Y}]$). Let $\pi\in\Pi(\mathrm{q},\mathrm{q}^{\prime})$ be a GW OT plan between $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$, i.e., it achieves $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}^{\prime},\mathrm{q}^{\prime}))=0$. Apply the blow-up procedure between $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$ with respect to $\pi$ (see Definition 4.2 and the follow up properties). Therefore, after relabeling the nodes if necessary, one obtains a new set of nodes $Y_{b}=Z=Y_{b}^{\prime}$, with cardinality equal to that of the support of $\pi$, and new distribution of mass $\mu_{Y_{b}}=\mu_{Z}=\mu_{Y_{b}^{\prime}}$ on such nodes. Moreover, there exist functions $\varphi_{Y}:Z\to Y$, $\varphi_{Y^{\prime}}:Z\to Y^{\prime}$ (defined implicitly in parts 2 and 3 of Definition 4.2) such that $(\varphi_{Y})_{\#}\mu_{Z}=\mu_{Y}$ $(\varphi_{Y^{\prime}})_{\#}\mu_{Z}=\mu_{Y^{\prime}}$ and $(\varphi_{Y})^{*}\omega_{Y}=(\varphi_{Y^{\prime}})^{*}\omega_{Y^{\prime}}$.
Now, consider $[v],[v^{\prime}]\in\mathrm{Tan}_{[\mathbb{Y}]}$ with representatives $v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})$ and $v^{\prime}\in L^{2}(Y^{\prime}\times Y^{\prime},\mu_{Y^{\prime}}\otimes\mu_{Y^{\prime}})$. As a consequence of Lemma B.11, we can obtain new representatives $(\varphi_{Y})^{*}v$ and $(\varphi_{Y^{\prime}})^{*}v^{\prime}$ defined over the same underlying space $Z\times Z$, i.e., $(\varphi_{Y})^{*}v,(\varphi_{Y^{\prime}})^{*}v^{\prime}\in L^{2}(Z\times Z,\mu_{Z}\otimes\mu_{Z})$, satisfying $(\varphi_{Y})^{*}v\simeq v$ and $(\varphi_{Y^{\prime}})^{*}v^{\prime}\simeq v$.
@@
Case $S=1$: For simplicity, we start by fixing only one finite template $[\mathbb{X}]$, and considering the functional $[\mathbb{Y}]\longmapsto\frac{1}{2}GW^{2}(\mathbb{X},\mathbb{Y})$ defined over classes of finite networks, which we denote by $G$. Given a finite network $[\mathbb{Y}]$, we want to compute $\nabla G([\mathbb{Y}])$ according to Definition 4.1.
Let $(\mathbf{X},\mathrm{p})$ and $(\mathbf{Y},\mathrm{q})$ be representatives of the weak classes of equivalence of $[\mathbb{X}]$ and $[\mathbb{Y}]$ after blow-up (for simplicity in the notation we avoid the subindex ‘$b$’ from Definition 4.2). In particular, both $\mathbf{X}$ and $\mathbf{Y}$ are matrices of the same size, $M\times M$. Moreover, we can assume that a GW OT plan in $\Pi(\mathrm{p},\mathrm{q})$ for $GW(\mathbb{X},\mathbb{Y})=GW((\mathbf{X},\mathrm{p}),(\mathbf{Y},\mathrm{q}))$ is induced by the identity map, after permuting and relabeling the nodes if necessary, and so we can also assume $\mathrm{p}=\mathrm{q}$. Thus,
$GW^{2}(\mathbb{X},\mathbb{Y})=\sum_{i,j\in[M]}|\mathbf{X}_{ij}-\mathbf{Y}_{ij}|^{2}\mathrm{q}_{i}\mathrm{q}_{j},$ (54)
@@
Consider the $M\times M$ matrix $\mathbf{X}-\mathbf{Y}$. The candidate for $\nabla G([\mathbb{Y}])$ is $[\mathbf{X}-\mathbf{Y}]$. Let $[v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ be an arbitrary tangent vector at $[\mathbb{Y}]$. We want to show that
@@
From Remark B.12 applied to $[v]$ and $[\mathbf{X}-\mathbf{Y}]$ in $\mathrm{Tan}_{[\mathbb{Y}]}$, we can consider representatives that are matrices of the same size. Therefore, without loss of generality, and by abuse of notation, we can assume that $v$ and $\mathbf{X}-\mathbf{Y}$ are representatives of $[v]$ and $[\mathbf{X}-\mathbf{Y}]$, respectively, of the same size $M\times M$ (after a new blow-up, possible enlargement, realignment, and relabeling). Thus,
$\langle[\mathbf{X}-\mathbf{Y}],[v]\rangle_{\mathrm{Tan}_{[\mathbb{Y}]}}=\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j},$ (55)
@@
For $t\geq 0$, consider $\pi^{t}$ a GW optimal $M\times M$ coupling in $\Pi(\mathrm{q},\mathrm{q})$ between $(\mathbf{X},\mathrm{q})$ (representative of $[\mathbb{X}]$) and $(\mathbf{Y}+tv,\mathrm{q})$ (representative of $Exp_{[\mathbb{Y}]}(tv)$) such that $\lim_{t\to 0}\pi^{t}=\mathrm{diag}(\mathrm{q})$ (notice that $Exp_{[\mathbb{Y}]}(0)=[\mathbb{Y}]$). On the one hand, we have:
@@
$\geq GW^{2}(\mathbb{X},\mathbb{Y})+t^{2}\sum_{i,j\in[M]}|v_{ij}|^{2}\mathrm{q}_{i}\mathrm{q}_{j}+2t\sum_{i,j\in[M]}\sum_{k,l\in[M]}(\mathbf{X}_{ik}-\mathbf{Y}_{jl})v_{jl}\pi_{ij}^{t}\pi^{t}_{kl}.$ (56)
@@
$=GW^{2}(\mathbb{X},\mathbb{Y})+t^{2}\sum_{i,j\in[M]}|v_{ij}|^{2}\mathrm{q}_{i}\mathrm{q}_{j}+2t\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j}.$ (57)
@@
    $\lim_{t\to 0}$ $\frac{GW^{2}(\mathbb{X},Exp_{[\mathbb{Y}]}(tv))-GW^{2}(\mathbb{X},\mathbb{Y})}{t}\leq 2\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j}.$ (58)
@@
    $\lim_{t\to 0}$ $\frac{GW^{2}(\mathbb{X},Exp_{[\mathbb{Y}]}(tv))-GW^{2}(\mathbb{X},\mathbb{Y})}{t}\geq 2\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j}.$ (59)
As a consequence of (55), (58) and (59), for a fixed $[\mathbb{X}]$, the directional derivative of the functional $G([\mathbb{Y}])=\frac{1}{2}GW^{2}(\mathbb{X},\mathbb{Y})$ in the direction of $[v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ at $[\mathbb{Y}]$ is given by
@@
Therefore, the gradient of the functional $[\mathbb{Y}]\longmapsto\frac{1}{2}GW^{2}(\mathbb{X},\mathbb{Y})$ can be identified with the matrix $\mathbf{X}-\mathbf{Y}$, obtained after blow-up.
@@
General Case $S\geq 1$: Let us fix a collection $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$ of classes of finite networks, $\lambda\in\Delta_{S-1}$, and the functional $G:\mathcal{GW}\to\mathbb{R}_{\geq 0}$ given in (31). Given a finite network $[\mathbb{Y}]$, we want to compute $\nabla G([\mathbb{Y}])$ according to Definition 4.1.
@@
Consider the $M\times M$ matrix $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}$. The candidate for $\nabla G([\mathbb{Y}])$ is $[\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}]\in\mathrm{Tan}_{[\mathbb{Y}]}$. Let $[v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ be an arbitrary tangent vector at $[\mathbb{Y}]$. As before, we can consider representatives of $[\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}]$ and $[v]$ that are matrices of the same size. By abuse of notation, we denote such representatives as $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}$ and $v$, assuming they belong to the space of $M\times M$ matrices endowed with the weighted Frobenius norm, with weight $\mathrm{q}\otimes\mathrm{q}$:
@@
$=\sum_{s\in[S]}\lambda_{s}\sum_{i,j\in[M]}\left(\mathbf{X}_{ij}^{s}-\mathbf{Y}_{ij}\right)v_{ij}\,\mathrm{q}_{i}\mathrm{q}_{j},$ (60)
@@
Let $[\mathbb{Y}]$ be any finite network class, and let us apply Algorithm 3 with respect to the given templates $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$ to obtain the corresponding blow-ups. On the one hand, we have that the blow-up 3-tuple $(Y_{b},\omega_{Y_{b}},\mu_{Y_{b}})$ is a representative of the class $\mathbb{[Y]}$. On the other hand, as a consequence of Proposition 4.3,
@@
[…]
@@
$\omega_{Y_{b}}=\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}^{s}}\quad\text{ or, equivalently, }\quad\mathbf{Y}_{b}=\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}.$ (61)
See also [14, Rmk. 9]. It is important to remark that the expression $\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}^{s}}$ depends on the given network $\mathbb{Y}$. That is, given another network $\mathbb{Z}$, one has to perform Algorithm 3 to $\mathbb{Z}$ and the templates $\{\mathbb{X}^{s}\}$ getting new representatives of all the networks involved and then the new corresponding expression $\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}}$ will be different to the above one (61). Thus, it would be better to denote the new representatives of the templates after blow-up with respect to $\mathbb{Y}$ as
@@
$[\mathbb{Y}^{*}]\in\mathrm{Bary}(\{[\mathbb{X}^{s}]\}_{s\in[S]})\Longrightarrow\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right).$
@@
We proceed as in the proof of Theorem 3.3 on the fixed-point approach. Since one can consider the identity map on $[M_{b}]$ as an optimal map for the problem $GW((\mathbf{X}^{s}_{b},\mathrm{q}_{b}),(\mathbf{Y}^{*}_{b},\mathrm{q}_{b}))$, in this case, the functional $J_{\lambda}:\mathbb{R}^{M_{b}\times M_{b}}\to\mathbb{R}_{\geq 0}$ can be written as
@@
This can be deduced from (38) or simply because minimizing $J_{\lambda}$ coincides with the barycentric problem within the Euclidean geometry (see (43) in App. A). Also, since $\mathbb{Y}^{*}$ is a GW barycenter, we have

### v2

@@ B.3. Weak Isomorphisms
Let $\mathbb{Y}=(Y,\omega_{Y},\mu_{Y})=(\mathbf{Y},\mathrm{q})$ and $\mathbb{Y}^{\prime}=(Y^{\prime},\omega_{Y^{\prime}},\mu_{Y^{\prime}})=(\mathbf{Y}^{\prime},\mathrm{q}^{\prime})$ be finite networks such that $\mathbb{Y}\sim^{w}\mathbb{Y}^{\prime}$ (i.e., both belong to the same class $\mathbb{Y}]$). Let $\pi\in\Pi(\mathrm{q},\mathrm{q}^{\prime})$ be a GW OT plan between $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$, i.e., it achieves $GW((\mathbf{Y},\mathrm{q}),(\mathbf{Y}^{\prime},\mathrm{q}^{\prime}))=0$. Apply the blow-up procedure between $\mathbb{Y}$ and $\mathbb{Y}^{\prime}$ with respect to $\pi$ (see Definition 4.2 and the follow up properties). Therefore, after relabeling the nodes if necessary, one obtains a new set of nodes $Y_{b}=Z=Y_{b}^{\prime}$, with cardinality equal to that of the support of $\pi$, and new distribution of mass $\mu_{Y_{b}}=\mu_{Z}=\mu_{Y_{b}^{\prime}}$ on such nodes. Moreover, there exist functions $\varphi_{Y}:Z\to Y$, $\varphi_{Y^{\prime}}:Z\to Y^{\prime}$ (defined implicitly in parts 2 and 3 of Definition 4.2) such that $(\varphi_{Y})_{\#}\mu_{Z}=\mu_{Y}$ $(\varphi_{Y^{\prime}})_{\#}\mu_{Z}=\mu_{Y^{\prime}}$ and $(\varphi_{Y})^{*}\omega_{Y}=(\varphi_{Y^{\prime}})^{*}\omega_{Y^{\prime}}$.
Now, consider $v],[v^{\prime}]\in\mathrm{Tan}_{[\mathbb{Y}]}$ with representatives $v\in L^{2}(Y\times Y,\mu_{Y}\otimes\mu_{Y})$ and $v^{\prime}\in L^{2}(Y^{\prime}\times Y^{\prime},\mu_{Y^{\prime}}\otimes\mu_{Y^{\prime}})$. As a consequence of Lemma B.11, we can obtain new representatives $(\varphi_{Y})^{*}v$ and $(\varphi_{Y^{\prime}})^{*}v^{\prime}$ defined over the same underlying space $Z\times Z$, i.e., $(\varphi_{Y})^{*}v,(\varphi_{Y^{\prime}})^{*}v^{\prime}\in L^{2}(Z\times Z,\mu_{Z}\otimes\mu_{Z})$, satisfying $(\varphi_{Y})^{*}v\simeq v$ and $(\varphi_{Y^{\prime}})^{*}v^{\prime}\simeq v$.
@@
Case $S=1$: For simplicity, we start by fixing only one finite template $\mathbb{X}]$, and considering the functional $\mathbb{Y}]\longmapsto\frac{1}{2}GW^{2}(\mathbb{X},\mathbb{Y})$ defined over classes of finite networks, which we denote by $G$. Given a finite network $\mathbb{Y}]$, we want to compute $\nabla G([\mathbb{Y}])$ according to Definition 4.1.
Let $(\mathbf{X},\mathrm{p})$ and $(\mathbf{Y},\mathrm{q})$ be representatives of the weak classes of equivalence of $\mathbb{X}]$ and $\mathbb{Y}]$ after blow-up (for simplicity in the notation we avoid the subindex ‘$b$’ from Definition 4.2). In particular, both $\mathbf{X}$ and $\mathbf{Y}$ are matrices of the same size, $M\times M$. Moreover, we can assume that a GW OT plan in $\Pi(\mathrm{p},\mathrm{q})$ for $GW(\mathbb{X},\mathbb{Y})=GW((\mathbf{X},\mathrm{p}),(\mathbf{Y},\mathrm{q}))$ is induced by the identity map, after permuting and relabeling the nodes if necessary, and so we can also assume $\mathrm{p}=\mathrm{q}$. Thus,
(54) $GW^{2}(\mathbb{X},\mathbb{Y})=\sum_{i,j\in[M]}|\mathbf{X}_{ij}-\mathbf{Y}_{ij}|^{2}\mathrm{q}_{i}\mathrm{q}_{j},$
@@
Consider the $M\times M$ matrix $\mathbf{X}-\mathbf{Y}$. The candidate for $\nabla G([\mathbb{Y}])$ is $\mathbf{X}-\mathbf{Y}]$. Let $v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ be an arbitrary tangent vector at $\mathbb{Y}]$. We want to show that
@@
From Remark B.12 applied to $[v]$ and $\mathbf{X}-\mathbf{Y}]$ in $\mathrm{Tan}_{[\mathbb{Y}]}$, we can consider representatives that are matrices of the same size. Therefore, without loss of generality, and by abuse of notation, we can assume that $v$ and $\mathbf{X}-\mathbf{Y}$ are representatives of $[v]$ and $\mathbf{X}-\mathbf{Y}]$, respectively, of the same size $M\times M$ (after a new blow-up, possible enlargement, realignment, and relabeling). Thus,
(55) $\langle[\mathbf{X}-\mathbf{Y}],[v]\rangle_{\mathrm{Tan}_{[\mathbb{Y}]}}=\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j},$
@@
For $t\geq 0$, consider $\pi^{t}$ a GW optimal $M\times M$ coupling in $\Pi(\mathrm{q},\mathrm{q})$ between $(\mathbf{X},\mathrm{q})$ (representative of $\mathbb{X}]$) and $(\mathbf{Y}+tv,\mathrm{q})$ (representative of $Exp_{[\mathbb{Y}]}(tv)$) such that $\lim_{t\to 0}\pi^{t}=\mathrm{diag}(\mathrm{q})$ (notice that $Exp_{[\mathbb{Y}]}(0)=[\mathbb{Y}]$). On the one hand, we have:
@@
(56) $\geq GW^{2}(\mathbb{X},\mathbb{Y})+t^{2}\sum_{i,j\in[M]}|v_{ij}|^{2}\mathrm{q}_{i}\mathrm{q}_{j}+2t\sum_{i,j\in[M]}\sum_{k,l\in[M]}(\mathbf{X}_{ik}-\mathbf{Y}_{jl})v_{jl}\pi_{ij}^{t}\pi^{t}_{kl}.$
@@
(57) $=GW^{2}(\mathbb{X},\mathbb{Y})+t^{2}\sum_{i,j\in[M]}|v_{ij}|^{2}\mathrm{q}_{i}\mathrm{q}_{j}+2t\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j}.$
@@
    (58) $\lim_{t\to 0}$ $\frac{GW^{2}(\mathbb{X},Exp_{[\mathbb{Y}]}(tv))-GW^{2}(\mathbb{X},\mathbb{Y})}{t}\leq 2\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j}.$
@@
    (59) $\lim_{t\to 0}$ $\frac{GW^{2}(\mathbb{X},Exp_{[\mathbb{Y}]}(tv))-GW^{2}(\mathbb{X},\mathbb{Y})}{t}\geq 2\sum_{i,j\in[M]}(\mathbf{X}_{ij}-\mathbf{Y}_{ij})v_{ij}\mathrm{q}_{i}\mathrm{q}_{j}.$
As a consequence of (55), (58) and (59), for a fixed $\mathbb{X}]$, the directional derivative of the functional $G([\mathbb{Y}])=\frac{1}{2}GW^{2}(\mathbb{X},\mathbb{Y})$ in the direction of $v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ at $\mathbb{Y}]$ is given by
@@
Therefore, the gradient of the functional $\mathbb{Y}]\longmapsto\frac{1}{2}GW^{2}(\mathbb{X},\mathbb{Y})$ can be identified with the matrix $\mathbf{X}-\mathbf{Y}$, obtained after blow-up.
@@
General Case $S\geq 1$: Let us fix a collection $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$ of classes of finite networks, $\lambda\in\Delta_{S-1}$, and the functional $G:\mathcal{GW}\to\mathbb{R}_{\geq 0}$ given in (30). Given a finite network $\mathbb{Y}]$, we want to compute $\nabla G([\mathbb{Y}])$ according to Definition 4.1.
@@
Consider the $M\times M$ matrix $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}$. The candidate for $\nabla G([\mathbb{Y}])$ is $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}]\in\mathrm{Tan}_{[\mathbb{Y}]}$. Let $v]\in\mathrm{Tan}_{[\mathbb{Y}]}$ be an arbitrary tangent vector at $\mathbb{Y}]$. As before, we can consider representatives of $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}]$ and $[v]$ that are matrices of the same size. By abuse of notation, we denote such representatives as $\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}-\mathbf{Y}$ and $v$, assuming they belong to the space of $M\times M$ matrices endowed with the weighted Frobenius norm, with weight $\mathrm{q}\otimes\mathrm{q}$:
@@
(60) $=\sum_{s\in[S]}\lambda_{s}\sum_{i,j\in[M]}\left(\mathbf{X}_{ij}^{s}-\mathbf{Y}_{ij}\right)v_{ij}\,\mathrm{q}_{i}\mathrm{q}_{j},$
@@
Let $\mathbb{Y}]$ be any finite network class, and let us apply Algorithm 3 with respect to the given templates $\{[\mathbb{X}^{s}]\}_{s=1}^{S}$ to obtain the corresponding blow-ups. On the one hand, we have that the blow-up 3-tuple $(Y_{b},\omega_{Y_{b}},\mu_{Y_{b}})$ is a representative of the class $\mathbb{[Y]}$. On the other hand, as a consequence of Proposition 4.3,
@@
$Exp_{[\mathbb{Y}]}([\nabla G_{\lambda}([\mathbb{Y}])])=[(Y_{b},\cancel{\omega_{Y_{b}}}+\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}^{s}}-\cancel{\omega_{Y_{b}}},\mu_{Y_{b}})]=[(Y_{b},\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}^{s}},\mu_{Y_{b}})].$
@@
(61) $\omega_{Y_{b}}=\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}^{s}}\quad\text{ or, equivalently, }\quad\mathbf{Y}_{b}=\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b}.$
See also [15, Rmk. 9]. It is important to remark that the expression $\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}^{s}}$ depends on the given network $\mathbb{Y}$. That is, given another network $\mathbb{Z}$, one has to perform Algorithm 3 to $\mathbb{Z}$ and the templates $\{\mathbb{X}^{s}\}$ getting new representatives of all the networks involved and then the new corresponding expression $\sum_{s=1}^{S}\lambda_{s}\omega_{X_{b}}$ will be different to the above one (61). Thus, it would be better to denote the new representatives of the templates after blow-up with respect to $\mathbb{Y}$ as
@@
$\mathbb{Y}^{*}]\in\mathrm{Bary}(\{[\mathbb{X}^{s}]\}_{s\in[S]})\Longrightarrow\mathbb{Y}^{*}\sim^{w}\left(\sum_{s=1}^{S}\lambda_{s}\mathbf{X}^{s}_{b},\mathrm{q}_{b}\right).$
@@
We proceed as in the proof of Theorem 3.3 on the fixed-point approach. Since one can consider the identity map on $M_{b}]$ as an optimal map for the problem $GW((\mathbf{X}^{s}_{b},\mathrm{q}_{b}),(\mathbf{Y}^{*}_{b},\mathrm{q}_{b}))$, in this case, the functional $J_{\lambda}:\mathbb{R}^{M_{b}\times M_{b}}\to\mathbb{R}_{\geq 0}$ can be written as
@@
This can be deduced from (37) or simply because minimizing $J_{\lambda}$ coincides with the barycentric problem within the Euclidean geometry (see (43) in Section A). Also, since $\mathbb{Y}^{*}$ is a GW barycenter, we have